Calculate the concentrations of the major ionic species present in the 1.29 mL solution of 3.22 mg of Ba(HCO3)2.
1
Expert's answer
2020-09-22T03:35:17-0400

Solution:

The molar mass of Ba(HCO3)2 is 259.36 g/mol.

Moles of Ba(HCO3)2 = Mass of Ba(HCO3)2 / Molar mass of Ba(HCO3)2

Moles of Ba(HCO3)2 = (0.00322 g) / (259.36 g/mol) = 0.0000124 mol


Molarity of Ba(HCO3)2 = Moles of Ba(HCO3)2 / Volume of solution

Molarity of Ba(HCO3)2 = (0.0000124 mol) / (0.00129 L) = 0.0096 mol/L = 0.0096 M


Dissociation reaction:

Ba(HCO3)2 ⟶ Ba2+ + 2HCO3-

The major ionic species: Ba2+; HCO3-


[Ba2+] = Molarity of Ba(HCO3)2 = 0.0096 M (according to the dissociation reaction)

[HCO3-] = 2 × Molarity of Ba(HCO3)2 = 2 × (0.0096 M) = 0.0192 M (according to the dissociation reaction)


Answer: The concentrations of the major ionic species: [Ba2+] = 0.0096 M; [HCO3-] = 0.0192 M.